### v2

# Large Language Models are Easily Confused: A Quantitative Metric, Security Implications and Typological Analysis

Abstract
Language Confusion is a phenomenon where Large Language Models (LLMs) generate text that is neither in the desired language, nor in a contextually appropriate language. This phenomenon presents a critical challenge in text generation by LLMs, often appearing as erratic and unpredictable behavior. We hypothesize that there are linguistic regularities to this inherent vulnerability in LLMs and shed light on patterns of language confusion across LLMs. We introduce a novel metric, Language Confusion Entropy, designed to directly measure and quantify this confusion, based on language distributions informed by linguistic typology and lexical variation. Comprehensive comparisons with the Language Confusion Benchmark  (Marchisio et al., 2024) confirm the effectiveness of our metric, revealing patterns of language confusion across LLMs. We further link language confusion to LLM security and find patterns in the case of multilingual embedding inversion attacks. Our analysis demonstrates that linguistic typology offers theoretically grounded interpretation, and valuable insights into leveraging language similarities as a prior for LLM alignment and security. [^1] The language graphs for language similarities and code are publicly available https://github.com/siebeniris/QuantifyingLanguageConfusion/
Large Language Models are Easily Confused:
A Quantitative Metric, Security Implications and Typological Analysis
Yiyi Chen1, QiongXiu Li2, Russa Biswas1, Johannes Bjerva1 1Department of Computer Science, 2Department of Electronic Systems Aalborg University, Copenhagen, Denmark {yiyic, rubi, jbjerva}@cs.aau.dk, qili@es.aau.dk
1 Introduction
Multilingual Large Language Models (LLMs) revolutionized Natural Language Processing (NLP), offering crosslinguality in various applications, including translation Zhu et al. (2024), text generation Chen et al. (2022), and information retrieval Guo et al. (2024). Besides the challenges faced by LLMs such as bias and fairness Talat et al. (2022), hallucinations Augenstein et al. (2024), multilingual LLMs are more vulnerable to adversarial and inversion attacks than monolingual LLMs Song et al. (2024); Chen et al. (2025, 2024).
Multilingual LLMs are trained on data in a diverse range of languages to represent the intricacies of multiple languages within a single model. However, this often results in inconsistencies in comprehension and response, leading to language confusion – instances where LLMs generate text that is neither in the desired language nor in a contextually appropriate one. For example, when an LLM is queried/prompted in Arabic, it may respond in text that is either partially or entirely in languages other than Arabic, e.g., English.
Marchisio et al. (2024) propose metrics to measure the percentage of model responses containing no undesired languages at both line and word levels but fail to capture nuances within language distributions. It is observed that language confusion tends to occur when the model’s distribution over the next tokens is flat. We hypothesize that language confusion in LLMs is not merely a performance limitation but an inherent vulnerability, partly due to imbalanced pre-training multilingual data sources, which are amplified with increasing numbers of languages and can be analyzed through language similarities derived from linguistic typology and other resources.
To thoroughly investigate language confusion as a phenomenon and its role within LLMs, we introduce the following research questions:
RQ1: What measurable patterns characterize language confusion in LLMs, and how can these patterns be quantified effectively?
RQ2: How do language similarities influence language confusion, and how can this knowledge be applied to enhance LLM alignment and security?
To this end, we propose a novel metric called Language Confusion Entropy, which provides a quantifiable measure of uncertainty and facilitates the detection of when an LLM is confused. Building on observations by Marchisio et al. (2024) that uniformity of the distribution indicates higher uncertainty, Language Confusion Entropy re-weights language distributions by emphasizing long-tail distributions, effectively capturing language confusion in multilingual LLM generation tasks. Furthermore, we demonstrate that this metric uncovers patterns of language confusion during both the training and evaluation phases of multilingual inversion attacks (Chen et al., 2025).
In addition, we construct language graphs based on language similarities derived from linguistic resources to analyze language confusion, revealing a strong correlation between language confusion and semantic similarities between languages. Our analysis shows that low-resource languages exhibit less confusion while training across diverse scripts and language families mitigates language confusion more effectively than training within the same script or language family in inversion attacks. These findings indicate that leveraging language similarities grounded in linguistic resources could serve as a valuable prior for enhancing LLM alignment and security. Our main contributions are as follows:
1) We propose a novel metric Language Confusion Entropy to measure language confusion in LLMs considering the nuances of language distributions. To the best of our knowledge, we are the first to quantify language confusion probabilistically.
2) Using language graphs, we demonstrate that linguistic typology provides a foundational tool for analyzing language confusion.
3) We propose a modified KL-Divergence algorithm to determine the correlation between language similarities (as defined by language graphs) and language confusion in LLMs.
4) We conduct extensive analysis revealing statistically significant patterns of language confusion, providing new insights for LLM security research.
2 Related Works
Language Confusion
This phenomenon observed in NLP, is often described as “off-target translation” (Chen et al., 2023a; Sennrich et al., 2024) or “accidental translation” (Zhang et al., 2020; Xue, 2020), or as “source language hallucinations” in zero-shot transfer scenarios (Vu et al., 2022; Li and Murray, 2023; Pfeiffer et al., 2023; Chirkova and Nikoulina, 2024). Language confusion, a term coined by Marchisio et al. (2024) occurs when the LLMs’ outputs are generated erroneously in languages different from the desired (target) languages and identified as ‘surprising limitation’ diminishing LLM utility for non-English languages, indicating the unpredictable nature.
The phenomenon of language confusion has not only been pervasive in LLMs, but also in tasks pertinent to LLM security, such as multilingual inversion attacks (Chen et al., 2024, 2025). Furthermore, Chen et al. (2025) observes language confusion across 20 languages from diverse scripts and language families in multilingual embedding inversion. They analyzed the pattern of confusion using basic typological features between train and eval languages with regression analysis, in comparison, the proposed Language Confusion Entropy provides a more interpretable analysis.
Multilingual LLM Safety and Security
Yong et al. (2024) exposes vulnerabilities of AI safety mechanism by jailbreaking GPT-4’s safeguard through translating unsafe English inputs into low-resource languages. Deng et al. (2024) impose unintentional and intentional jailbreak on multilingual LLMs, using multilingual prompts. It is observed that low-resource languages are more vulnerable, making them the weakest links in AI security.
Backdoor attacks on multilingual machine translation pose significant threats, as injecting poisoned data into low-resource language pairs can achieve a high attack success rate (ASR) in high-resource language pairs (Wang et al., 2024). Poisoning instruction-tuning data for one or two languages can affect other languages, surpassing 99% ASR in the cross-lingual setting in prominent LLMs resisting current defenses (He et al., 2024).
Multilingual textual embedding inversion attacks pose additional risks, as any encoder can be attacked to reconstruct original texts. Traditional defenses for monolingual LLMs are ineffective for multilingual LLMs (Chen et al., 2024, 2025). Moreover, Song et al. (2024) generates language blending for adversarial attacks, necessitating systematic analysis of language similarity and language confusion for targeted defenses.
Linguistic Typology and Language Similarities
Previous research on multilingual effects on linguistic level uses three approaches: (i) phylogenetic variation, (ii) linguistic typological variation, and (iii) embedded and data-driven language variation.
While genealogical relations are intuitive, the correlations between language similarity and genealogical relations are often spurious (Rama and Kolachina, 2012). Ploeger et al. (2024) challenge this approach highlighting its negative impact on downstream NLP tasks.
Linguistic typology offers a theoretically grounded approach to measuring similarity between languages (Kashyap, 2019). Languages can be categorized based on various features — e.g., whether they use a Subject-Verb word order (SV) or the opposite (VS). Such information has been manually annotated in linguistic databases, including WALS (Haspelmath, 2008), ASJP  (Wichmann et al., 2012), and Grambank  (Skirgård et al., 2023), and work in NLP has contributed with automatic prediction of such features (Malaviya et al., 2017; Bjerva et al., 2019a; Bjerva, 2024). Recent work has also explored lexically driven measures, showing that multilingual LLMs often rely on lexical overlap (Pires et al., 2019). Such work spans from using synonymy-relations as in WordNet (Fellbaum, 2010), to multilingual relations in BabelNet (Navigli and Ponzetto, 2010), and more complex colexification patterns in CLICS3 (Rzymski et al., 2020). Crosslingual colexification patterns refer to the phenomenon whereby different meanings are captured by the same lexica across languages (François, 2008), implicating shared cognitive or cultural associations (Karjus et al., 2021; Di Natale et al., 2021; Chen and Bjerva, 2023).
Language embeddings encode language characteristics and can be derived from data-driven methods or linguistic typological databases. Östling and Tiedemann (2017) trained a character-level LSTM language model on translated Bible texts from 990 languages, showing their ability to reconstruct language genealogies. Additionally, embeddings can be generated from typological data sources like WALS, Grambank, and ASJP. Chen et al. (2023b) created embeddings using lexical data based on colexification patterns in CLICS3 and BabelNet.
We hypothesize that language confusion is a phenomenon largely driven by lexical variation, similar to the patterns observed by Pires et al. (2019). We investigate this by building our analysis on this body of work leveraging computational typology.
3 Explainable Language Confusion
To investigate the phenomenon of language confusion, we use the datasets i) Language Confusion Benchmark (LCB) (Marchisio et al., 2024) and ii) Multilingual Textual Embedding Inversion (MTEI) (Chen et al., 2025) (see Appendix for task details and Table 8 for processed datasets sample).
Figure 1: The use of proposed metric to quantify language confusion and its correlation with language similarity through KL divergence.
3.1 Generation Settings
Consider a LLM trained or prompted with a set of $n\in\mathbb{N}$ source languages $L_{s}=\{l_{1},\cdots,l_{n}\}$, and $l_{t}$ is the target language. We probe language confusion for both LLM instruction and textual embedding inversion attacks.
Monolingual Generation
LCB: The model is queried in language $l_{t}$, and the response is expected in $l_{t}$. MTEI: The inversion model, trained on $l_{t}$, inverts embeddings in $l_{t}$. Here, $l_{t}\in L_{s}$, i.e., the evaluated language is part of the training languages.
Crosslingual Generation
LCB: The model is instructed in language $l_{s}$ to provide a response in $l_{t}$, where $l_{t}\neq l_{s}$. MTEI: The inversion model, trained on $L_{s}$, inverts embeddings in $l_{t}$. In this setting, $l_{t}\notin L_{s}$, meaning the evaluated language differs from the training languages.
3.2 Quantifying Language Confusion
The phenomenon of language confusion is particularly prominent in crosslingual generation settings. For instance in Fig. 1  1, when an LLM is prompted in English and expected to generate a response in German ($X_{1}$), the output may unexpectedly have a mix of other languages, such as Spanish and French ($X_{2}$). Ideally, the LLM should focus on the expected languages; thus, the model exhibits greater confusion if its output distribution assigns high probabilities to unexpected languages. To quantify this, we propose Language Confusion Entropy ($H_{\mathbf{C}}$), defined as follows:
$H_{\mathbf{C}}(X)=$ $-\sum_{x\in X_{1}}(1-p(x))\log(p(x))$ (1)
$-\sum_{x\in X_{2}}p(x)\log p(x),$
where $X_{1}$ denotes the expected language set and $X_{2}$ the unexpected language set, $X_{1}\cup X_{2}=X,~{}X_{1}\cap X_{2}=\emptyset$, $p(x)$ denotes the probability and $\sum_{x\in X}p(x)=1$. Ideally, an unconfused model would satisfy $\sum_{x\in X_{1}}p(x)=1$ and $\sum_{x\in X_{2}}p(x)=0$.
In practical applications, for the LCB dataset, $X_{1}$ includes both the instruction and response languages, while for the MTEI dataset, $X_{1}$ encompasses the training and evaluation languages.
3.3 Language Identification
Language confusion can occur at both the word level and the line level. To measure it, we use existing language identification (LID) tools. Initially, we employ Lingua to detect languages as it offers the highest accuracy among existing LID tools, particularly at word level. [^2] https://github.com/pemistahl/lingua-py However, since Lingua supports only 75 languages, we supplement it with fastText (Joulin et al., 2016) (which supports 176 languages) for languages unidentified by Lingua.
Line-level Detection
We split the generated output into lines by newline characters and detect the language of each line.
Word-level Detection
To detect languages more accurately at the word level, we first tokenize the text at the line level using language-specific tokenizers such as jieba (Sun et al., 2013) (Chinese), Hebrew Tokenizer (Levin and Oriyan, 2018) Kiwipiepy (Lee, 2024) (Korean), fugashi (McCann, 2020) (Japanese), and NLTK Word Tokenizer  (Bird and Loper, 2004) for other languages. We then identify the language of each tokenized word.
Finally, we compile the detected languages into distributions at both word and line levels for further analysis. Examples of the pre-processed data are presented in Table 8. Language confusion matrices are then constructed by applying language confusion entropy to these distributions, and the results are aggregated per language (ref. Fig. 1  1).
Algorithm 1 KL Divergence for Language Confusion vs. Language Similarity Matrices
[…]
2:Mean KL divergence across all columns.
[…]
4:for each column index $j$ from $1$ to $m$ do
[…]
7:     Step 1: Exclude zeros from M1_col
[…]
11:     Step 2: Normalize the distributions
12:     $P\leftarrow P/\sum P$
13:     $Q\leftarrow Q/\sum Q$
14:     Avoid division by zero or log issues
15:     $\epsilon\leftarrow 10^{-10}$
16:     $P\leftarrow P+\epsilon$
17:     $Q\leftarrow Q+\epsilon$
18:     Step 3: Calculate KL divergence
[…]
20:     Append KL_Div to Total_KL_Divergence
21:end for
[…]
3.4 Large Language Models
The evaluated LLMs in LCB include Command R (35B parameters),Command R+ (104B), GPT-3.5 Turbo (Brown, 2020), and GPT-4 Turbo (Achiam et al., 2023), Mistral Large, Mistral 8x7B (Jiang et al., 2024), LLaMA 2 70B Instruct (Touvron et al., 2023), and LLaMA 3 70B Instruct, while in  MTEI the inversion models are trained with mT5 (580M) (Xue, 2020) and multilingual-e5-base (me5) (580M). (See Table 5 for details of LLMs).
3.5 Language Graphs
We construct language graphs from a diverse range of typological features, such as colexification patterns (Rzymski et al., 2020; Fellbaum, 2010; Navigli and Ponzetto, 2010), lexicon (Wichmann et al., 2012), phonological and morphological-syntactical features (Haspelmath, 2008; Skirgård et al., 2023), as well as from a collection of existing language embeddings trained from NLP tasks incorporating linguistic typology (Östling and Tiedemann, 2017; Östling and Kurfalı, 2023; Chen et al., 2023b). We then generate language similarity matrices from the language graphs by calculating pairwise similarity using either Jaccard Index or Cosine Similarity (ref. Fig. 1  2). (See details in Appendix B).
3.6 The Role of Language Similarity in Language Confusion
We compare the language graphs with language confusion matrices to assess how well language confusion aligns with language similarities and to identify specific aspects where they match. To quantify the divergence between language confusion (denoted as $P$) and language similarity (denoted as $Q$), we employ Kullback-Leibler Divergence (Kullback and Leibler, 1951), expressed as $KL(P||Q)$. $P(x)$ represents the distribution of language confusion entropy of language $x$ relative to other languages, while $Q(x)$ represents the distribution of language similarity of $x$ relative to other languages. $KL(P||Q)$ is computed as follows:
$KL(P||Q)=\sum_{x}P(x)log\left(\frac{P(x)}{Q(x)}\right),$ (2)
where a lower $KL(P||Q)$ indicates a stronger correspondence between language confusion patterns and underlying language similarities. (See Algorithm 1 for detailed steps and Fig 1  3).
Figure 2: Language confusion for LCB by each language across LLMs for crosslingual setting at Line level. The languages are ordered ascendingly by their language confusion entropy averaged across LLMs.
4 Analysis and Results
4.1 Language Confusion in LLM Prompting
Language Confusion Entropy vs. Pass Rates
The binary metrics, Pass Rates at line-level LPR and word-level WPR are used to evaluate whether the LLM output contains no error, following Marchisio et al. (2024) (see details in Appendix A). We apply language confusion entropy to LCB, calculating it at both the line-level $H_{\mathbf{C}}$[L] and word-level $H_{\mathbf{C}}$[W] across generation settings.
Compared to Marchisio et al. (2024), our approach detects language confusion across all languages, including at the word level, by using language-specific tokenizers and a more accurate language identification (LID) tool. We reproduce LPR and WPR (ref. Table 1011) and compute $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] (ref. Table 9) in both crosslingual and monolingual settings, for 14 languages and 8 LLMs, following (Marchisio et al., 2024).
To evaluate the efficacy of language confusion entropy compared to pass rate metrics, we calculate the Spearman correlation [^3] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.spearmanr.html coefficients between these metrics across levels and generation settings. Overall, $H_{\mathbf{C}}$[L] shows a strong negative correlation with LPR across all generation settings. Moreover, $H_{\mathbf{C}}$[W] - which is based on more detailed language distributions - exhibits a weaker correlation with WPR, as WPR only considers English words in non-Latin script languages. Despite the weaker correlation, it remains statistically significant for all languages, especially crosslingually.
At both the line and word levels, $H_{\mathbf{C}}$ shows a stronger correlation with pass rates in crosslingual settings than in monolingual ones. This aligns with the definition of language confusion entropy, which gives more weight to long-tail distributions, a more prominent phenomenon in crosslingual tasks.
$H_{\mathbf{C}}$[L] $H_{\mathbf{C}}$[W] LPR
All $H_{\mathbf{C}}$[W] $0.51$***
LPR $-0.83$*** $-0.3$***
WPR $-0.29$** $\mathbf{-0.5}$*** $0.31$**
Monolingual $H_{\mathbf{C}}$[W] $0.42$***
LPR $-0.72$*** $-0.23$*
WPR $0.01$ $-0.31$ $0.08$
Crosslingual $H_{\mathbf{C}}$[W] $0.54$***
LPR $\mathbf{-0.87}$*** $-0.38$***
WPR $-0.27$** $-0.47$*** $\mathbf{0.37}$***
Table 1: Spearman correlation between Language Confusion Entropy and Pass Rates at both Word level and Line level with LCB. The strongest correlation is in bold.
Figure 3: Language confusion entropy for LCB across generation settings by LLMs at line and word Level.
Figure 4: Language confusion for LCB across data sources at line level for crosslingual setting.
Figure 5: Language Confusion for LCB across generation settings by data sources at line and word Level.
Language Confusion Entropy Across LLMs
As shown in Table 9 and Fig. 3, language confusion is more likely to occur in crosslingual compared to monolingual, with each LLM presenting significant variance. Word-level language confusion presents more variance per LLM and higher severity than line-level. Overall, it is consistent that Command and GPT LLMs have relatively lower language confusion than Mistral and LLaMA LLMs, projecting similar findings from Marchisio et al. (2024).
There is a clear consistency in language confusion across different LLMs, particularly at the line level and in crosslingual settings, as shown in Fig. 3. LLaMA 3 70B-I consistently exhibits the highest confusion across nearly all languages, while GPT-4 Turbo demonstrates the lowest confusion, especially for high-resource languages like French, Spanish, German, and Chinese. Command R+ also shows relatively low confusion across most languages, except Indonesian.
Notably, languages that are written in non-Latin scripts (on the right side of the X-axis), such as Vietnamese, Chinese, Korean, Arabic, Japanese, Hindi, and Indonesian, consistently show higher confusion entropies across most LLMs, especially in the LLaMA models. In contrast, Latin-script languages like French, Spanish, German, and Italian tend to have lower confusion rates across all LLMs, particularly in GPT-4 Turbo and Command R+.
Language Confusion Entropy Across Data Sources
The data for monolingual and crosslingual tasks in LCB consists of 2,600 and 4,500 prompts, respectively, sourced from 7 datasets (details in Table 4). As shown in Fig. 4, language confusion is more pronounced in crosslingual settings and at the word level. Monolingually, language confusion tends to align with the median word length (W) of prompts in each dataset. For example, Aya, Dolly, Okapi, and Native prompts have median lengths of 9, 10, 13, and 19 words, respectively, and their language confusion follows this order. Crosslingually, the Complex Prompts dataset has the highest median word length (159, compared to 18 for ShareGPT and 15 for Okapi), and it also exhibits the highest language confusion.
Observing language confusion across datasets at the line level for crosslingual settings (Fig. 5), [^4] We use “Language Confusion Entropy” and “Confusion Entropy” interchangeably in this paper. a clear pattern emerges. Complex Prompts has the highest confusion across all languages, while ShareGPT shows the lowest confusion for most languages, except for French and Spanish. Consistent with previous findings, languages written in non-Latin scripts show higher confusion in datasets like Okapi and ShareGPT. However, in Complex Prompts, non-Latin-script languages such as Chinese, Korean, Arabic, and Japanese demonstrate lower confusion than Latin-script languages.
Figure 6: Language Confusion and Text Reconstruction Performance (BLEU) in Multilingual Textual Embedding Inversion Attacks at Line Level for Crosslingual Settings for Eval (Top) and Train (Bottom) Languages.
4.2 Language Confusion in Multilingual Textual Embedding Inversion Security
Language Confusion Entropy for Eval Languages When embeddings are in languages that are more likely to be confused, they are more prone to being inverted into text in "incorrect" languages, reducing the inversion performance, especially with word-matching metrics like BLEU (Post, 2018). Also, the languages generated by the inversion model are often skewed by the pre-training data of the LLM, such as mT5 in MTEI.
$H_{\mathbf{C}}$[L] $H_{\mathbf{C}}$[W] BLEU
All $H_{\mathbf{C}}$[W] $0.89$***
BLEU $\mathbf{-0.62}$*** $\mathbf{-0.44}$***
mT5 $0.71$*** $0.62$*** $-0.32$*
Monolingual $H_{\mathbf{C}}$[W] $0.75$***
BLEU $0.25$ $0.17$
mT5 0.1 0.26 $-0.59$***
Crosslingual $H_{\mathbf{C}}$[W] $\mathbf{0.9}$***
BLEU $-0.48$*** $-0.36$**
mT5 $\mathbf{0.76}$*** $\mathbf{0.66}$*** $-0.32$*
Table 2: Spearman Correlations among $H_{\mathbf{C}}$ at Line Level and Word Level and BLEU score for MTEI, and the percentage of pre-training data in mT5 for eval languages. Strongest correlations are in bold.
To test this intuition, we calculate the Spearman correlation among language confusion entropy, BLEU scores, and the percentage of respective languages in the pre-training data of mT5 (see Table 7 for details). As shown in Table 2, for eval languages, $H_{\mathbf{C}}$[L] is strongly correlated with inversion performance across generation settings, confirming that language confusion negatively impacts reconstruction performance.
Moreover, $H_{\mathbf{C}}$[L] and $H_{\mathbf{C}}$[W] are both strongly correlated with the proportion of languages in pre-training data of mT5, particularly in crosslingual setting. This indicates that languages with higher representation in the pre-training data are more prone to confusion, both at the word and line levels. This observation is further validated by empirical evidence: when an inversion model trained on Arabic is used to invert embeddings in Meadow Mari (unseen in mT5), the model often generates English, highlighting the influence of pre-training data on embedding inversion attacks. Fig 6 (top) provides a visualization of language confusion at the line level for each language, alongside their corresponding reconstruction performance in BLEU at each step of the evaluation in the crosslingual inversion attacks. It shows directly that lower-resourced languages present lower confusion, especially for non-Latin script languages, whereas they are also more vulnerable in terms of higher reconstruction performance, for example, Gujarati, Punjabi, and Urdu.
Language Graph LCB Textual Embedding Inversion
AVG ALL Monolingual Crosslingual AVG ALL Monolingual Crosslingual
Line Word Line Word Line Word Line Word Line Word Line Word
Grambank 0.2650 0.1830 0.2827 0.2726 0.4042 0.1791 0.2682 0.7538 0.8005 0.7891 0.7504 0.5434 0.8230 0.8162
WALS 0.1947 0.0420 0.2908 0.0816 0.4286 0.0388 0.2865 0.7377 0.8722 0.7854 0.6618 0.4102 0.8803 0.8164
[…]
Lang2Vec
Inventory 0.1650 0.0812 0.2003 0.1374 0.3114 0.0637 0.1961 0.8154 0.9678 0.8410 0.8225 0.4124 0.9720 0.8768
Syntactic 0.2925 0.1949 0.3771 0.2244 0.4679 0.1505 0.3405 0.8651 1.0179 0.8668 0.8872 0.5447 1.0001 0.8742
Phonological 0.2260 0.1009 0.3126 0.1552 0.4235 0.0830 0.2808 1.3161 1.7178 1.6133 0.6859 0.6004 1.6599 1.6191
Genetic 12.8307 13.4548 12.2278 12.9350 12.3563 13.5850 12.4249 14.9443 14.6988 14.9756 18.5227 13.5734 13.8721 14.0232
ASJP SVD 0.5859 0.0253 1.0597 0.0522 1.0395 0.0225 1.3164 2.8254 2.9730 3.3461 0.0820 1.6568 3.8882 5.0063
ASJP UMAP 0.9318 0.0994 1.6767 0.1290 1.6055 0.0946 1.9856 3.3217 3.5364 4.0245 0.0811 2.3479 4.3690 5.5711
Colex2Lang
CLICS 0.1665 0.0289 0.2522 0.0589 0.3776 0.0260 0.2490 0.7333 0.9492 0.8335 0.3690 0.2672 0.9503 0.8693
WN 0.1489 0.0242 0.2154 0.0522 0.3404 0.0218 0.2149 0.7794 1.0105 0.8892 0.3263 0.3894 1.0445 0.9454
WN_CONCEPT 0.1490 0.0242 0.2175 0.0522 0.3426 0.0218 0.2168 0.7791 1.0100 0.8836 0.3263 0.3894 1.0427 0.9390
Table 3: KL Divergence between Language Similarity Graphs and the Language Confusion Matrices for Target/ Eval Languages from LCB and Inversion Tasks. The best results (lowest) are bolded, and the second best are underlined.
Language Confusion Entropy for Train Languages
In MTEI, the inversion models are trained in three different settings - monolingual, in-family and in-script, and cross-family and cross-script. As shown in Fig. 6 (bottom) and Table 13, monolingual training renders lower language confusion for each train language while pairing training languages, in-script/in-family training renders higher language confusion compared to cross-script/cross-family training. These findings substantiate the intuition that similar languages are more prone to confusion.
Our study reveals that inversion performance significantly improves when trained in in-script/in-family settings (ref. Table 13 in the Appendix). Crosslingual inversion performances are comparable to in-script/in-family training when trained in Kazakh (Latin-script) combined with Gujarati and Punjabi, respectively, and language confusion is notably lower. This suggests that while similar languages tend to increase confusion, certain crosslingual combinations can achieve strong performance without the added confusion seen in in-family/in-script training. Overall, these findings highlight the trade-off between inversion performance and language confusion, indicating further optimization is needed to strike the ideal balance between them.
4.3 Language Confusion and Linguistic Typology
Table 3 shows the best results from KL divergence between language confusion and language similarities based on different language graphs for both LCB and MTEI, using Algorithm 1, the whole results are presented in Table 15 in Appendix.
Our findings reveal strong correlations between language confusion and language similarities based on various typological sources. For instance, the similarity measures based on semantic typology correlate the most strongly, followed closely by more general lexical similarity measures. Language similarities based on typological feature databases like Grambank and WALS show stronger correlations than those based on parallel Bible texts (Östling and Tiedemann, 2017). Interestingly, and echoing previous findings on typological variation, we find genetic variation is a poor proxy for this analysis (Bjerva et al., 2019b; Ploeger et al., 2024), indicating the need for theoretically grounded approaches to linguistic interpretation.
5 Discussion
Language Confusion Entropy $H_{\mathbf{C}}$ vs. Pass Rate (PR)
As shown in Section 4.1, $H_{\mathbf{C}}$ and PR are correlated but measure distinct aspects of language generation. Importantly, $H_{\mathbf{C}}$ is not intended to replace PR but to offer a deeper quantitative characterization of language confusion.
$H_{\mathbf{C}}$ quantifies behavioral uncertainty through the entropy of a model’s probability distribution across languages. It captures partial confusions (e.g., 40% confidence in the target language, 60% spread across others), even when outputs technically “pass.” PR is a binary performance metric assessing whether outputs meet target language requirements.
Hence, $H_{\mathbf{C}}$ complements PR by revealing hidden confusion patterns: a model with high PR could still exhibit high $H_{\mathbf{C}}$ if probabilities scatter across non-target languages. This makes $H_{\mathbf{C}}$ particularly valuable for diagnosing how models arrive at correct outputs, not just whether they do.
Research Questions
In response to RQ1, we proposed an effective metric Language Confusion Entropy, through which we identified several patterns contributing to language confusion. These include prompt complexity, imbalanced distributions of training sources, and language similarities - all play a significant role in language confusion. Furthermore, our findings indicate that these factors strongly correlate with inversion performance and the pretraining languages in LLMs.
Inherent vulnerabilities in LLMs stem from intrinsic design, training processes, or model architectures, which are not directly caused by attacks or improper use. However, external attacks can amplify and leverage the vulnerabilities, which are directly reflected in embedding inversion attacks. Further, decoding unexpected languages disrupts the user experience. We have discussed that i) language confusion stems from training data inequality, ii) the phenomena is pervasive from related works, iii) language confusion is well-captured by the proposed LCE and shows consistent trends across different architectures of LLMs and datasets and experimental settings, and (iv) can be influenced and explained by linguistic typological variations. These affords a positive response to RQ2.
Moreover, there is a potential to leverage language similarities as a prior for LLM alignment and security. For instance, Table 3 shows that the language similarities from colexification patterns afford a strong correlation with language confusion (with low KL divergence), which indicates that LLMs easily confuse languages that contain words that are crosslingually capturing the same senses. When an LLM is exposed to multilingual data with more distinct colexification patterns, it could enhance its ability to distinguish them and make them more resilient against language confusion. This strategy could promote more resilient LLMs, as we have shown that models are less likely to confuse typologically dissimilar languages. Hence, exploring typology-aware design strategies could provide both offensive and defensive insights in LLM security.
Potential Misuse
The Language confusion metric identifies uncertainty in language identification, highlighting areas where LLMs may be prone to errors. Our findings show that language similarities correlate with language confusion patterns in LLMs. When similar languages lack robust safety measures compared to well-protected high-resource languages, they could be exploited for crosslingual attacks in a more targeted manner, such as backdoors and jailbreaking (Zou et al., 2023; Li et al., 2024). Additionally, the ability of LLMs to switch between languages may pose risks where safety measures aren’t consistently implemented across all languages. The Language confusion metric identifies uncertainty in language identification, highlighting areas where LLMs may be prone to errors. Our findings show that language similarities correlate with language confusion patterns in LLMs. When similar languages lack robust safety measures compared to well-protected high-resource languages, they could be exploited for crosslingual attacks in a more targeted manner, such as jailbreaking (He et al., 2024) and backdoors (Wang et al., 2024). Additionally, the ability of LLMs to switch between languages may pose risks where safety measures aren’t consistently implemented across all languages. It has been demonstrated that LLMs are particularly vulnerable to crosslingual attacks in related work (Wang et al., 2024; He et al., 2024) and more recently (Poppi et al., 2024), mainly because the conventional defense mechanisms designed for monolingual settings are ineffective in multilingual settings. To enhance LLM security, adversarial exploits can be detected by monitoring high language confusion entropy, especially in crosslingual settings. Our work suggests that language confusion quantification and its connection to language similarities can be leveraged to raise awareness of such vulnerabilities, and also provide potential revenue for developing mitigation strategies.
6 Conclusion and Future Work
Addressing the challenge of language confusion, we introduce Language Confusion Entropy, a novel metric that quantifies language confusion by re-weighting language distributions and emphasizing long-tail patterns. This metric captures language confusion in multilingual LLM tasks, revealing patterns of uncertainty in both training and evaluation phases. Our findings show strong correlations between language confusion and semantic similarities among languages, with less confusion observed in low-resource languages and when training incorporates diverse scripts and language families. These insights confirm that language confusion fundamentally impacts LLMs and suggest linguistic typology as a potential tool for enhancing model security. Detecting language confusion enables smoother and more precise interactions in multilingual contexts while enhancing user trust in LLM-based AI systems across domains such as legal services and healthcare. In future work, we aim to apply these findings to practical applications, such as developing typology-aware defense to improve LLM alignment and security. Key applications include cross-lingual chatbots, translation services, detecting code-switching, and improving multilingual speech recognition by reducing ambiguity.
Limitations
A core limitation of this work is that some of our analysis and downstream implications can only be carried out on languages that are represented in typological databases. When part of the work, limited to typological databases, inspires downstream solutions in terms of defense mechanisms, undocumented languages may not benefit from these advances. However, we have also increased coverage of languages in constructing language graphs using data-driven methods. Our core method is also not limited to any typological database.
Ethics Statement
This work adheres to the ACL ethics guidelines. We investigate language confusion and link findings to security vulnerabilities of low-resource languages, including those using non-Latin scripts and with diverse typologies. The potential misuse has been extensively discussed. Our work highlights how these factors can be used to potentially improve the security of low-resource language technology. We encourage the community to incorporate a broader range of languages in NLP security research, to ensure that low-resource languages are also covered by defense mechanisms developed in the future.
Acknowledgements
YC and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454).
Appendix A Datasets for Language Confusion
Dataset name Nature of Data $|L|$ $|D|$ Languages $W$
Monolingual Aya (Iyer et al., 2024) Human-generated 100 500 eng, tur, arb, cmn, por 9
Dolly (Iyer et al., 2024) MT post-edited 100 500 hin, rus, fra, arb, esp 10
Okapi (Lai et al., 2023) Synthetic+MT 100 1.2k eng, fra, ita, deu, cmn, vie, rus, esp, find, por, arb, hin, esp, fra, jap, kor 13
Native prompts (Marchisio et al., 2024) Human-generated 100 500 esp, fra, jap, kor 19
Crosslingual Okapi (Lai et al., 2023) Synthetic 100 1.5k $\mathcal{L}$ 15
ShareGPT (https://sharegpt.com/) Human-generated 100 1.5k $\mathcal{L}$ 18
Complex prompts (Marchisio et al., 2024) Human-generated 99 1.5k $\mathcal{L}$ 159
Table 4: Data Sources in the LCB for monolingual and crosslingual generation (Marchisio et al., 2024). $|D|$ is the total number of examples per data source and $|L|$ is the number of examples per language. For the crosslingual setting, the model is instructed in English to generate in the target language $l\in\mathcal{L}$ where $\mathcal{L}$=fra, deu, esp, por, ita, jap, kor, cmn, arb, tur, hin, rus, ind, vie. $W$ is the median length in words of the prompts in each dataset.
Language Confusion Benchmark
Marchisio et al. (2024) create and release a language confusion benchmark covering 15 languages, sourcing prompts from publicly available multilingual instruction datasets, and also creating new data with more complex promts, see detailed data sources in Table 4.
Additionally, the binary metrics such as Line Pass Rate (LPR) and Word Pass Rate (WPR) are defined in Marchisio et al. (2024) to measure whether a response contains any instance of a) a line in an incorrect language and b) an isolated English word/phrase for languages using non-Latin scripts.
LPR calculates the percentage of model responses that pass the line-level language confusion detector without error. A response is “correct” if all lines match the user’s desired language.
$LPR=\frac{|R\backslash E_{L}|}{|R|}$ (3)
where $R$ is the set of all responses and $E_{L}$ is the set of responses that contain line-level errors.
WPR measures the percentage of responses where all words are in the desired language.
$WPR=\frac{|(R\backslash E_{L})|\backslash E_{W}}{|R\backslash E_{L}|}$ (4)
where $R$ is the set of all responses, $E_{L}$ is the set of responses with line-level errors, and $E_{W}$ the set of responses with word-level errors.
We reproduce the LPR and WPR on LCB for both crosslingual and monolingual settings (as shown in Table 10 and 11). The detailed results applying language confusion entropy to LCB are presented in Table 9, for comparison.
LLM Transformer #Languages Parameters Reference
LCB
Command R Decoder-only - 35B https://cohere.com/blog/command-r
Command R+ Decoder-only - 104B https://cohere.com/blog/command-r-plus-microsoft-azure
GPT-3.5 Turbo Decoder-only - - Brown (2020)
GPT-4 Turbo Decoder-only - - Achiam et al. (2023)
Mistral Large Decoder-only - - https://mistral.ai/news/mistral-large/
Mistral 8x7B Decoder-only - 7B Jiang et al. (2024),
Llama 2 70B Instruct Decoder-only - 70B Touvron et al. (2023)
Llama 3 70B Instruct Decoder-only - 70B https://ai.meta.com/blog/meta-llama-3/
MTEI
mT5-base Encoder-Decoder 102 580M Xue (2020)
multilingual-e5-base Encoder 94 580M https://huggingface.co/intfloat/multilingual-e5-base
Table 5: The Evaluated LLMs.
Multilingual Textual Embedding Inversion
Textual embedding inversion has presented a standing challenge in LLM security, where the private texts can be reconstructed from evasdroped embeddings from Embeddings as a Service (EaaS), by training an attacker model based on the embeddings extracted from the black-box embedders (Song and Raghunathan, 2020; Lyu et al., 2020; Kim et al., 2022; Morris et al., 2023; Chen et al., 2024, 2025). However, most work was done in monolingual settings, mostly in English, other than the recent work expands the language space to four Romance and Germanic languages in Latin script  (Chen et al., 2024) and in Chen et al. (2025), the inversion atacks are extended to 20 languages across 8 families and 12 scripts (see Table 7). The trained inversion attack model consists of a base model and a corrector model, where a base model is a text-to-text generation model, while a corrector model is used to bring closer the generated embeddings and attacked embeddings in the embedding space. While in the evaluation phase, three stages are reported: base, step1 (corrector model) and step50+sbeam8 (corrector model with beam search with sequence length 8). The inversion model is trained with mT5 (Xue, 2020) as base model and multilingual-e5-base  [^5] Huggingface: intfloat/multilingual-e5-base as black-box encoder. The samples of the curated dataset are shown in Table 8.
We apply language confusion entropy to eval and train languages in the monolingual and crosslingual settings at both line and word levels, while comparing the BLEU score in the regarding scenario (ref. Table 12 and 13).
$H_{\mathbf{C}}$[L] $H_{\mathbf{C}}$[W] BLEU
All $H_{\mathbf{C}}$[W] $0.93$***
BLEU $0.66$*** $0.71$***
mT5 $0.32$*** $0.24$ $0.09$
Monolingual $H_{\mathbf{C}}$[W] $0.63$***
BLEU $0.37$*** $0.33$**
mT5 0.05 $0.42$* $0.09$
Crosslingual $H_{\mathbf{C}}$[W] $0.93$***
BLEU $0.66$*** $0.71$***
mT5 $0.32$ $0.23$ $0.09$
Table 6: Spearman Correlations among $H_{\mathbf{C}}$ at Line Level and Word Level and BLEU score for Multilingual Textual Embedding Inversion, and the percentage of pre-training data in mT5, for train languages.
Language ISO 639 Lang. Family Lang. Script Script ISO Directionality #Samples(Train) WO mT5 (%)
Arabic arb Semitic Arabic Arab RTL 1M VSO 1.66
Urdu urd Indo-Aryan Arabic Arab RTL 600K SOV 0.61
Kazakh kaz Turkic Cyrillic Cyrl LTR 1M SOV 0.65
Mongolian mon Mongolic Cyrillic Cyrl LTR 1M SOV 0.62
Hindi hin Indo-Aryan Devanagari Deva LTR 600K SOV 1.21
Gujarati guj Indo-Aryan Gujarati Gujr LTR 600K SOV 0.43
Punjabi pan Indo-Aryan Gurmukhi Guru LTR 600K SOV 0.37
Chinese cmn Sino-Tibetan Haqniqdoq Hani LTR 1M SVO 1.67
Hebrew heb Semitic Hebrew Hebrewr RTL 1M SVO 1.06
Japanese jpn Japonic Japanese Jpan LTR 1M SOV 1.92
German deu Germanic Latin Latn LTR 1M Non-Dominant 3.05
Turkish tur Turkic Latin Latn LTR 1M SOV 1.93
Amharic amh Semitic Ethiopian Ethi LTR - SOV 0.29
Sinhala sin Indo-Aryan Sinhala Sinh LTR - SOV 0.41
Korean kor Koreanic Hangul Hang LTR - SOV 1.14
Finnish fin Uralic Latin Latn LTR - SVO -
Hungarian hun Uralic Latin Latn LTR - Non-Dominant 1.48
Yiddish ydd Germanic Hebrew Hebrewr RTL - SVO 0.28
Maltese mlt Semitic Latin Latn LTR - Non-Dominant 0.64
Meadow Mari mhr Uralic Cyrillic Cyrl LTR - SOV -
Table 7: Languages and their Language Characteristics, i.e., Language Family, Language Script, Directionality of the Script, Number of Training Samples for Inversion Models, Word Order of Subject, Object and Verb in Multilingual Inversion Attack (Chen et al., 2025) and the Percentage of the language in Pre-training data in mT5 (Xue, 2020).
Model $\mathcal{L}_{\text{train}}$ $\mathcal{L}_{\text{eval}}$ Eval Step Predicted Language Distribution
me5 Hindi German Base {eng: 0.27, deu: 0.34, hin: 0.24, fra: 0.01, nld: 0.01, fin: 0.01, mar: 0.02, nep: 0.01}
me5 Hindi German Step1 {eng: 0.17, deu: 0.47, hin: 0.24, mar: 0.02, fra: 0.01, nep: 0.01, nld: 0.01}
me5 Hindi German Step50+sbeam8 {hin: 0.38, mar: 0.04, deu: 0.37, eng: 0.15, fra: 0.01}
me5 Hindi Yiddish Base {mar: 0.12, hin: 0.83, eng: 0.02, deu: 0.01, nep: 0.01}
me5 Hindi Yiddish Step1 {hin: 0.82, eng: 0.02, mar: 0.13, deu: 0.01}
me5 Hindi Yiddish Step50+sbeam8 {hin: 0.81, mar: 0.12, deu: 0.01, eng: 0.04}
me5 Hindi Hebrew Base {hin: 0.92, eng: 0.03, mar: 0.03}
me5 Hindi Hebrew Step1 {hin: 0.94, mar: 0.03, eng: 0.02}
me5 Hindi Hebrew Step50+sbeam8 {hin: 0.94, eng: 0.02, mar: 0.03}
me5 Hindi Arabic Base {eng: 0.63, hin: 0.32, mar: 0.02, nep: 0.01, fra: 0.01}
me5 Hindi Arabic Step1 {eng: 0.61, hin: 0.33, mar: 0.03}
me5 Hindi Arabic Step50+sbeam8 {eng: 0.62, hin: 0.31, mar: 0.04}
Table 8: Examples of Dataset MTEI. The probabilities for unidentified languages are omitted.
AVG French Spanish Italian German Russian Portuguese Turkish Vietnamese Chinese Korean Arabic Japanese Hindi Indonesian
Monolingual (Line)
Command R 0.0306 0.0272 0.0381 0.0092 0.0430 0.0093 0.0195 0.0042 0.0099 0 0 0.0017 0.0171 0.0111 0.2384
Command R+ 0.0355 0.0127 0.0325 0 0.0373 0.0309 0.0219 0.0100 0.0264 0 0 0.0034 0.0158 0.0387 0.2673
GPT-3.5 Turbo 0.0317 0.0250 0.0296 0.0103 0.0276 0.0168 0.0292 0.0050 0.0070 0 0 0 0.0057 0.0546 0.2337
GPT-4 Turbo 0.0381 0.0508 0.0468 0.0051 0.0316 0.0636 0.0228 0 0.0033 0.0079 0 0 0.0083 0.0639 0.2289
Mistral 8x7B 0.0766 0.0399 0.0510 0.0645 0.0229 0.0414 0.0784 0.0252 0.1037 0.1116 0.0787 0.0583 0.1394 0.0540 0.2036
Mistral Large 0.0799 0.0145 0.0351 0.0202 0.0139 0.0156 0.0643 0.0378 0.2245 0.1011 0.0329 0.1231 0.0756 0.0663 0.2936
Llama 2 70B-I 0.1266 0.0729 0.0548 0.1500 0.1438 0.0704 0.0569 0.1347 0.2501 0.1113 0.0501 0.1713 0.0784 0.1279 0.2999
Llama 3 70B-I 0.1197 0.0483 0.0557 0.0300 0.0556 0.0955 0.0654 0.2009 0.1416 0.1518 0.1720 0.1879 0.1579 0.1720 0.1412
Crosslingual (Line)
Command R 0.1555 0.1942 0.1651 0.1723 0.1928 0.1357 0.1458 0.1441 0.1579 0.1007 0.1169 0.1159 0.1522 0.1395 0.2437
Command R+ 0.1216 0.1234 0.1635 0.0810 0.1040 0.1034 0.1555 0.0817 0.1204 0.0610 0.0617 0.0534 0.0691 0.1424 0.3825
GPT-3.5 Turbo 0.1025 0.1219 0.1085 0.1042 0.1119 0.0749 0.1169 0.0945 0.0772 0.0527 0.0596 0.0764 0.0624 0.1163 0.2584
GPT-4 Turbo 0.0987 0.0721 0.0956 0.1002 0.0778 0.0876 0.0909 0.0733 0.0742 0.0665 0.0600 0.0666 0.0981 0.1612 0.2577
Mistral 8x7B 0.1675 0.1341 0.1332 0.1242 0.1582 0.1304 0.1288 0.1156 0.1458 0.1604 0.1743 0.1506 0.1977 0.3053 0.2861
Mistral Large 0.2274 0.1341 0.2122 0.1808 0.1441 0.1496 0.2509 0.1410 0.1817 0.2944 0.3052 0.2816 0.3416 0.2407 0.3262
Llama 2 70B-I 0.2649 0.1517 0.1339 0.2330 0.2657 0.2186 0.2022 0.2280 0.2532 0.2817 0.3574 0.3719 0.3233 0.3443 0.3433
Llama 3 70B-I 0.4631 0.3906 0.3192 0.4324 0.4226 0.4495 0.3503 0.5696 0.5360 0.3872 0.3777 0.5200 0.4822 0.5531 0.6935
Monolingual (Word)
Command R 0.5563 0.9016 0.9424 0.9513 0.4098 0.6312 1.0095 0.5445 0.4059 0.1974 0.0031 0.1424 0.0462 0.4161 1.1865
Command R+ 0.5522 0.9034 0.9687 0.9363 0.3873 0.6478 1.0185 0.5003 0.3812 0.1232 0.0057 0.1327 0.0785 0.4007 1.2458
GPT-3.5 Turbo 0.5344 0.9031 0.9348 0.9567 0.4325 0.6322 0.9457 0.4281 0.3587 0.0913 0.0088 0.1053 0.0370 0.3937 1.2532
GPT-4 Turbo 0.5426 0.8656 0.9456 0.9284 0.4241 0.6206 0.9818 0.4932 0.3217 0.1287 0.0146 0.1355 0.0307 0.4232 1.2823
Mistral 8x7B 0.7431 0.9763 0.9157 1.0027 0.4477 0.6947 0.9746 0.6763 0.4134 0.8854 0.4960 0.4330 0.8628 0.6106 1.0134
Mistral Large 0.5703 0.8829 0.9099 0.9083 0.4270 0.6191 0.9827 0.6473 0.3199 0.2191 0.2568 0.3137 0.2484 0.4472 0.8016
Llama 2 70B-I 0.8022 1.0541 0.9257 1.0109 0.8092 0.6939 1.0000 1.2333 0.5748 0.5989 0.3960 0.5578 0.5229 0.5676 1.2862
Llama 3 70B-I 0.8657 1.0519 0.9427 1.0139 0.8992 0.6325 1.0028 1.5459 0.4827 0.8949 0.8120 0.6904 0.6017 0.6831 0.8659
Crosslingual (Word)
Command R 0.6796 1.0941 1.0038 1.0896 0.6588 0.6618 0.9487 0.5579 0.4664 0.3347 0.1988 0.3190 0.4943 0.4836 1.2030
Command R+ 0.6049 0.9606 0.9380 0.9444 0.4382 0.6684 1.0785 0.5316 0.4572 0.2081 0.0670 0.1616 0.1870 0.5074 1.3212
GPT-3.5 Turbo 0.5951 1.0295 0.9439 0.9894 0.5858 0.5796 1.0159 0.5327 0.3659 0.1273 0.0933 0.1727 0.1850 0.4478 1.2626
GPT-4 Turbo 0.6348 1.0025 0.9619 0.9890 0.5257 0.6425 1.0182 0.5568 0.4089 0.2542 0.1182 0.2552 0.3012 0.5073 1.3451
Mistral 8x7B 0.8556 1.0509 1.0205 1.0143 0.6203 0.7656 1.0186 0.6174 0.5099 0.7446 0.6496 0.5800 0.9520 1.1017 1.3330
Mistral Large 0.9647 1.2351 1.1284 1.1409 0.6319 0.7850 1.1244 0.6909 0.6404 1.1388 0.7054 0.7411 1.2626 0.8266 1.4551
Llama 2 70B-I 0.9676 1.2830 0.9488 1.1583 0.9822 0.7885 1.0626 0.6962 0.6398 1.1089 0.8069 0.6929 1.1530 0.9268 1.2983
Llama 3 70B-I 0.9757 1.2111 1.1366 1.3193 0.8329 1.0911 1.2465 0.9568 0.7462 0.7249 0.3435 0.8228 0.8275 0.8733 1.5274
Table 9: Language Confusion measured by Language Confusion Entropy for LCB for monolingual and crosslingual settings at line and word level for each Language for LLMs.
LPR AVG French Spanish Italian German Russian Portuguese Turkish Vietnamese Chinese Korean Arabic Japanese Hindi Indonesian
Monolingual
Command R 98.50 99.33 95.67 99.00 98.00 100.00 98.50 99.00 99.00 98.50 100.00 100.00 100.00 100.00 92.00
Command R+ 99.19 99.67 99.33 100.00 100.00 100.00 97.50 100.00 99.00 97.50 100.00 99.67 99.00 100.00 97.00
GPT-3.5 Turbo 99.05 100.00 99.67 100.00 100.00 100.00 98.00 100.00 99.00 97.00 100.00 100.00 98.00 99.00 96.00
GPT-4 Turbo 99.26 99.33 99.33 99.00 100.00 100.00 98.00 100.00 100.00 99.00 100.00 99.00 100.00 100.00 96.00
Mistral 8x7B 71.08 95.33 89.33 72.00 91.00 65.00 85.00 90.00 57.00 45.50 61.00 48.00 67.00 71.00 58.00
Mistral Large 67.82 100.00 99.00 99.00 98.00 98.00 79.50 71.00 29.00 66.00 64.00 48.00 48.00 19.00 31.00
Llama 2 70B-I 44.65 87.67 95.67 72.00 59.00 89.00 91.00 33.00 17.00 10.50 0.00 0.33 7.00 1.00 62.00
Llama 3 70B-I 42.15 88.67 98.33 88.00 31.00 77.00 95.50 18.00 10.00 8.00 0.00 21.67 10.00 23.00 21.00
Crosslingual
Command R 77.84 86.83 84.17 74.00 72.00 77.00 79.80 75.50 74.00 84.40 77.00 80.83 74.00 74.25 76.00
Command R+ 93.77 95.50 95.50 95.25 93.75 94.75 92.00 94.00 93.00 92.80 93.25 96.50 95.00 92.75 88.75
GPT-3.5 Turbo 92.83 94.00 96.50 93.50 92.75 93.75 93.20 92.00 93.50 90.60 92.75 95.33 90.75 93.75 87.25
GPT-4 Turbo 93.13 95.00 96.17 93.50 94.75 92.50 93.80 93.50 92.50 92.40 92.25 94.00 90.75 93.25 89.50
Mistral 8x7B 69.73 87.17 84.17 81.75 80.00 70.50 81.60 79.50 71.00 52.60 58.00 53.50 60.25 47.25 69.00
Mistral Large 62.26 86.00 83.83 74.25 80.50 72.00 70.60 67.25 48.25 54.20 46.75 42.00 45.25 48.50 52.25
Llama 2 70B-I 40.53 79.33 86.50 67.75 54.00 51.00 82.00 26.25 19.55 10.84 3.58 6.33 14.00 16.25 50.00
Llama 3 70B-I 34.81 70.83 79.67 49.25 33.75 48.00 70.80 17.53 16.53 5.80 0.58 26.33 3.53 40.50 24.25
Table 10: Language Confusion Benchmark Line Pass Rate Reproduction for both Monolingual and Crosslingual settings.
WPR AVG Russian Chinese Korean Arabic Japanese Hindi
Monolingual
Command R 96.31 96.00 92.50 97.00 99.33 94.00 99.00
Command R+ 99.44 98.00 100.00 100.00 99.67 99.00 100.00
GPT-3.5 Turbo 99.83 100.00 100.00 100.00 100.00 99.00 100.00
GPT-4 Turbo 99.67 100.00 99.00 99.00 100.00 100.00 100.00
Mistral Large 98.50 99.00 99.00 100.00 100.00 98.00 95.00
Mistral 8x7B 73.75 83.00 64.50 62.00 86.00 68.00 79.00
Llama 2 70B-I 97.92 93.00 94.50 100.00 100.00 100.00 100.00
Llama 3 70B-I 93.19 94.00 93.50 100.00 95.67 80.00 96.00
Crosslingual
Command R 93.89 93.67 91.00 97.33 94.33 88.33 98.67
Command R+ 95.11 89.67 95.67 96.00 98.00 95.33 96.00
GPT-3.5 Turbo 98.72 98.33 99.00 98.33 99.00 98.67 99.00
GPT-4 Turbo 96.61 95.67 97.00 96.67 97.33 95.67 97.33
Mistral Large 93.83 93.00 97.67 92.00 93.67 91.33 95.33
Mistral 8x7B 68.22 80.33 62.00 67.33 76.33 51.67 71.67
Llama 2 70B-I 84.17 81.33 76.00 86.67 91.33 84.33 85.33
Llama 3 70B-I 94.39 89.00 92.33 100.00 95.67 91.67 97.67
Table 11: Language Confusion Benchmark Word Pass Rate Reproduction for both Monolingual and Crosslingual settings.
Language Step Monolingual Crosslingual BLEU (AVG)
Line Level Word Level Line Level Word Level
Sinhala Base - - 0.9061 1.6668 6.6147
Step1 - - 1.1412 1.7610 6.1822
Step50+sbeam8 - - 1.3213 1.9353 6.1517
Gujarati Base 0.0385 0.2332 0.9372 1.7629 8.8752
Step1 0.0308 0.9894 1.4092 1.7977 8.7153
Step50+sbeam8 0.0154 0.2303 1.3821 1.5844 10.0989
Meadow Mari Base - - 1.1895 2.5281 6.8674
Step1 - - 1.3452 2.4527 6.3181
Step50+sbeam8 - - 1.4333 2.4806 6.0744
Punjabi Base 0 0.2304 1.2886 1.9215 11.7362
Step1 0 0.1910 1.0879 1.7473 9.7027
Step50+sbeam8 0 0.1994 0.8049 1.3955 11.4941
Amharic Base - - 1.2949 1.9906 7.2613
Step1 - - 1.1420 1.7983 6.7393
Step50+sbeam8 - - 1.2548 1.8958 6.5141
Urdu Base 0.0516 0.7021 1.3266 2.3594 8.6339
Step1 0.0462 0.5510 1.4870 2.8171 8.2143
Step50+sbeam8 0.0385 0.5661 1.5844 2.3360 8.9373
Korean Base - - 1.4917 2.5477 5.5578
Step1 - - 1.6311 2.2823 5.1378
Step50+sbeam8 - - 1.6845 2.4643 4.9561
Yiddish Base - - 1.5216 2.3321 6.5486
Step1 - - 1.6912 2.2014 6.0670
Step50+sbeam8 - - 1.5956 2.2891 5.9480
Mongolian Base 0.0462 2.6470 1.6223 2.6724 7.9523
Step1 0.0462 2.6291 1.3172 2.5401 7.7916
Step50+sbeam8 0.0462 2.6586 1.4105 2.6977 8.4952
Hindi Base 0.6587 1.0295 2.0050 3.0441 8.5273
Step1 0.6305 1.0192 1.9432 2.5054 8.5912
Step50+sbeam8 0.6212 1.0268 1.4697 2.6635 9.6811
Arabic Base 0.0280 0.5021 2.1931 3.0200 6.1455
Step1 0 0.4597 1.7083 2.3915 5.9620
Step50+sbeam8 0 0.4666 1.6406 2.1459 6.1903
Kazakh Base 0.1390 1.5820 2.2796 3.1694 10.8145
Step1 0.1070 1.4129 1.8063 3.0509 10.8421
Step50+sbeam8 0.1098 1.4378 1.8065 3.0765 12.4864
Hebrew Base 0 0.1555 2.8174 3.7918 5.9312
Step1 0 0.1625 2.3715 3.3317 5.6492
Step50+sbeam8 0 0.1391 2.0314 2.5126 5.7413
Chinese Base 0 0.6124 2.8718 3.2475 6.4620
Step1 0.0280 0.5898 1.9708 2.3969 5.6923
Step50+sbeam8 0.0280 0.6262 1.9363 2.2672 5.3919
Maltese Base - - 3.0059 4.1352 6.7777
Step1 - - 3.5456 4.5662 6.1049
Step50+sbeam8 - - 3.3048 4.5243 5.8991
Japanese Base 0.0693 0.8806 3.2545 4.9242 5.7113
Step1 0.0693 0.8135 2.6345 3.2156 5.2014
Step50+sbeam8 0.0693 0.8277 2.3426 3.1882 5.0411
German Base 0.0000 0.6205 3.2764 4.6278 6.3678
Step1 0.0231 0.6106 3.8041 4.4014 6.3865
Step50+sbeam8 0 0.6203 3.3912 4.4915 6.8334
Turkish Base 0.0660 1.6099 3.8939 4.8367 7.9789
Step1 0.0707 1.6103 2.7842 3.7629 8.3830
Step50+sbeam8 0.0687 1.6045 2.8521 4.0266 9.6997
Finnish Base - - 4.3696 5.6543 5.2877
Step1 - - 3.6533 5.0938 4.8876
Step50+sbeam8 - - 3.6939 5.0801 4.7593
Hungarian Base - - 4.4119 5.1123 5.1684
Step1 - - 3.6595 4.6260 4.7418
Step50+sbeam8 - - 3.3812 4.0739 4.6179
Table 12: Language Confusion Entropy at each step in the Monolingual and Crosslingual generation settings at both Line and Word level, with BLEU score for textual embedding inversion tasks for eval languages.
Train Language Step Monolingual Crosslingual BLEU (AVG)
Line Level Word Level Line Level Word Level
Monolingual Punjabi Base 0.0000 0.2615 0.3042 0.7821 6.7660
Step1 0.0000 0.0000 0.0000 0.0000 0.0227
Step50+sbeam8 0.0000 0.0000 0.0000 0.0000 0.0229
Hebrew Base 0.0000 0.1720 0.3404 0.8445 6.0112
Step1 0.0000 0.1391 0.4184 0.8806 5.9759
Step50+sbeam8 0.0000 0.1391 0.5401 0.9759 5.9736
Gujarati Base 0.0462 0.2332 0.7272 1.6964 6.3759
Step1 0.0462 0.2384 0.6853 1.3227 6.4278
Step50+sbeam8 0.0462 0.2384 0.4426 1.0568 6.4171
Chinese Base 0.0000 0.5667 0.7711 1.3158 6.3859
Step1 0.0000 0.5667 0.6766 1.5313 6.1478
Step50+sbeam8 0.0560 0.6395 0.8930 1.2632 5.9179
German Base 0.0000 0.6205 0.8030 1.4859 6.5072
Step1 0.0000 0.6008 0.9991 1.4383 6.4497
Step50+sbeam8 0.0000 0.6090 1.0879 1.4596 6.4801
Kazakh Base 0.1252 0.9288 1.2115 2.1219 7.5891
Step1 0.1061 0.9288 0.7061 1.7686 7.5050
Step50+sbeam8 0.1252 0.9288 0.7562 1.7288 8.0075
Urdu Base 0.0462 0.7023 1.2381 1.9326 6.0338
Step1 0.0462 0.6597 0.9932 1.8497 5.9788
Step50+sbeam8 0.0462 0.6781 1.0471 1.8990 5.8903
Japanese Base 0.0925 0.8399 1.2677 2.5338 5.8317
Step1 0.0925 0.8265 1.3099 2.3773 5.6813
Step50+sbeam8 0.0925 0.8112 1.4349 2.5009 5.4885
Turkish Base 0.0925 1.0902 1.2699 2.1591 7.1899
Step1 0.0462 1.0711 0.9757 1.8074 7.2374
Step50+sbeam8 0.0462 1.0711 0.9824 2.0698 7.5047
Hindi Base 0.6952 1.0268 1.7053 2.0748 6.4369
Step1 0.6276 0.9830 1.5004 1.8830 6.3956
Step50+sbeam8 0.6417 1.0094 1.5470 1.9180 6.3725
Arabic Base 0.0560 0.5021 1.8446 2.1560 6.0938
Step1 0.0000 0.4471 0.8492 1.6087 6.0149
Step50+sbeam8 0.0000 0.4610 1.1300 1.3489 6.1260
Mongolian Base 0.0462 0.7688 2.2537 2.9952 7.0101
Step1 0.0462 0.7330 1.8297 2.9202 7.3608
Step50+sbeam8 0.0462 0.7697 1.6677 2.5683 7.8355
In-Family/In-Script Hindi, Punjabi Base 0.3333 0.6300 2.6597 3.4826 8.1018
Step1 0.3317 0.6300 3.3548 3.7614 8.5398
Step50+sbeam8 0.3209 0.6300 3.4952 4.2414 9.2197
Punjabi, Urdu Base 0.0393 0.4470 2.9478 3.9370 8.0232
Step1 0.0231 0.4361 4.0750 4.8806 7.3287
Step50+sbeam8 0.0231 0.4546 3.7656 4.6257 8.0605
Kazakh, Turkish Base 0.0995 1.0037 3.0674 5.2358 8.3012
Step1 0.1158 0.9838 2.3818 4.7775 8.7420
Step50+sbeam8 0.0924 1.0029 2.6656 4.2396 9.8136
Gujarati, Hindi Base 0.3209 0.6300 3.1485 4.7015 7.3518
Step1 0.3332 0.6389 3.2087 3.5149 7.5794
Step50+sbeam8 0.3369 0.6257 2.8926 3.1977 8.1357
German, Turkish Base 0.0231 2.7236 3.3850 5.2398 7.5745
Step1 0.0693 2.7123 3.3267 5.2520 7.9995
Step50+sbeam8 0.0393 2.7291 3.0071 5.3435 8.9119
Gujarati, Urdu Base 0.0462 0.4768 3.6012 4.2957 7.5971
Step1 0.0231 0.4678 3.6518 4.7169 7.8532
Step50+sbeam8 0.0231 0.4579 3.2620 4.0107 8.3644
Hindi, Urdu Base 0.3327 0.8645 3.6261 4.0579 7.4791
Step1 0.3258 0.8347 3.5240 3.8894 7.7025
Step50+sbeam8 0.3064 0.8460 3.2558 3.7994 8.1493
Arabic, Hebrew Base 0.0000 0.3206 3.9135 4.6422 7.2555
Step1 0.0000 0.3290 3.0216 3.8845 7.1863
Step50+sbeam8 0.0000 0.3056 3.0272 4.1192 7.3948
Gujarati, Punjabi Base 0.0231 0.2332 4.3682 5.0570 7.9168
Step1 0.0231 2.4993 3.1955 4.0486 8.2486
Step50+sbeam8 0.0000 0.2220 2.7857 4.2355 9.1628
Kazakh, Mongolian Base 0.0995 4.9836 4.6250 5.5072 9.0396
Step1 0.0761 4.9836 4.3483 5.1664 9.6267
Step50+sbeam8 0.0995 4.9947 3.6838 5.6001 10.8565
Cross-Family/Cross-Script Gujarati, Turkish Base 0.0693 0.6534 1.7507 3.1341 7.3780
Step1 0.0463 0.6525 2.4953 3.1636 7.6174
Step50+sbeam8 0.0463 0.6409 2.1876 3.1276 8.1549
Chinese, Japanese Base 0.0231 0.7897 1.9550 3.2433 6.4919
Step1 0.0511 0.7067 3.2448 4.7459 6.5438
Step50+sbeam8 0.0231 0.7285 3.7689 4.0670 6.6168
Punjabi, Turkish Base 0.0231 0.6417 2.1019 3.4377 8.0628
Step1 0.0231 0.6440 2.8690 3.6396 7.2908
Step50+sbeam8 0.0231 0.6548 2.5556 3.6179 7.9988
Hindi, Turkish Base 0.3682 1.0384 2.3767 3.5239 -
Step1 0.3489 1.0369 2.2821 3.1379 -
Step50+sbeam8 0.3601 1.0321 2.3311 3.1449 -
Hindi, Kazakh Base 0.4053 0.9859 2.8022 4.7000 -
Step1 0.4069 0.9749 3.2070 4.3596 -
Step50+sbeam8 0.3735 0.9994 3.4030 3.9206 -
Gujarati, Kazakh Base 0.0761 0.5911 2.8922 3.9972 7.8211
Step1 0.0857 0.5548 2.4678 3.0565 8.2088
Step50+sbeam8 0.0531 0.5810 2.0942 2.9441 8.8596
Turkish, Urdu Base 0.0462 0.9011 2.9759 4.0922 7.5098
Step1 0.0693 0.8723 2.3225 3.3931 7.7308
Step50+sbeam8 0.0462 0.9011 2.1042 3.3795 8.1401
Kazakh, Punjabi Base 0.0764 0.5697 3.2837 4.0113 8.4029
Step1 0.0531 0.5435 2.2224 3.2103 8.6699
Step50+sbeam8 0.0764 0.5705 1.6113 3.1262 9.3574
Table 13: Language Confusion Entropy at each step in the Monolingual and Crosslingual generation settings at both Line and Word level, with BLEU score for textual embedding inversion tasks for train languages.
Appendix B Language Graphs for Language Similarities
We curate language graphs from a diverse range of sources, as shown in Table 14. The language vectors from Grambank and WALS consist of multi-valued features, while those derived from colexification patterns in CLICS3 and WordNet (WN) are binarized. For these, we employ the Jaccard index to compute pairwise language similarities. For other more dense valued language vectors, we use cosine similarity instead.
Language Graph #Languages #Features Category of Features Datasource References Representation Similarity Metric
Grambank 2292 195 Morpho-syntactical Skirgård et al. (2023) Multivalued Vectors Jaccard
WALS 424 192 Structural properties of languages  Haspelmath (2008) Multivalued Vectors Jaccard
[…]
Lang2Vec (Littell et al., 2017) 8070 - Inventory, Syntax, Phonology, Genealogy, Geography, Featural  Collin (2010); Haspelmath (2008); Collins and Kayne (2011) Vectors Cosine similarity and Arccosine
$\text{CLICS}^{3}$ 1347 4228 Colexifications  Rzymski et al. (2020) Binarized Vectors Jaccard
WN 519 2,518,357 Colexifications  Navigli and Ponzetto (2010) Binarized Vectors Jaccard
Colex2Lang (Chen et al., 2023b)
$\text{CLICS}^{3}$ 1609 4228 Colexifications  Rzymski et al. (2020) Combined Graph Embeddings Cosine Similarity
WN 519 2,525,591 Colexifications  Navigli and Ponzetto (2010) Combined Graph Embeddings Cosine Similarity
WN_CONCEPT 519 2,486,485 Colexifications  Navigli and Ponzetto (2010) Combined Graph Embeddings Cosine Similarity
ASJP SVD 1012 40 lexicon  (Wichmann et al., 2012) UMAP on a mean normalized Levenshtein distance pairwise distance matrix from ASJP Cosine Similarity
ASJP UMAP 1012 40 lexicon  (Wichmann et al., 2012) Truncated SVD on a mean normalized Levenshtein distance pairwise distance matrix from word alignments Cosine Similarity
Östling and Tiedemann (2017) 943 - - Bible Translations Lang. Vectors trained on Bible Data Cosine Similarity
Table 14: Statistics and Incorporated Features for Language Graphs.
Language Graph LCB Textual Embedding Inversion
AVG ALL Monolingual Crosslingual AVG ALL Monolingual Crosslingual
Line Word Line Word Line Word Line Word Line Word Line Word
Grambank 0.2650 0.1830 0.2827 0.2726 0.4042 0.1791 0.2682 0.7538 0.8005 0.7891 0.7504 0.5434 0.8230 0.8162
WALS 0.1947 0.0420 0.2908 0.0816 0.4286 0.0388 0.2865 0.7377 0.8722 0.7854 0.6618 0.4102 0.8803 0.8164
[…]
Lang2Vec
Inventory 0.1650 0.0812 0.2003 0.1374 0.3114 0.0637 0.1961 0.8154 0.9678 0.8410 0.8225 0.4124 0.9720 0.8768
Syntactic 0.2925 0.1949 0.3771 0.2244 0.4679 0.1505 0.3405 0.8651 1.0179 0.8668 0.8872 0.5447 1.0001 0.8742
Phonological 0.2260 0.1009 0.3126 0.1552 0.4235 0.0830 0.2808 1.3161 1.7178 1.6133 0.6859 0.6004 1.6599 1.6191
Genetic 12.8307 13.4548 12.2278 12.9350 12.3563 13.5850 12.4249 14.9443 14.6988 14.9756 18.5227 13.5734 13.8721 14.0232
Geographical 1.5976 1.5924 1.8769 0.9753 1.8614 1.4996 1.7798 2.5218 2.4910 2.3964 2.6389 2.7875 2.4732 2.3439
Featural 0.3174 0.2094 0.4099 0.2429 0.5033 0.1626 0.3762 1.0628 1.0940 0.9355 1.5467 0.7976 1.0673 0.9359
$\text{CLICS}^{3}$ 1.6777 1.6516 1.7336 1.5453 1.7043 1.7182 1.7130 1.1806 1.1601 1.0850 1.4595 1.2880 1.0670 1.0237
WN 1.1372 1.1211 1.1423 1.1491 1.2521 1.0906 1.0680 2.1364 2.2197 2.0398 2.9402 1.6996 2.0687 1.8501
Colex2Lang
$\text{CLICS}^{3}$ 0.1665 0.0289 0.2522 0.0589 0.3776 0.0260 0.2490 0.7333 0.9492 0.8335 0.3690 0.2672 0.9503 0.8693
WN 0.1489 0.0242 0.2154 0.0522 0.3404 0.0218 0.2149 0.7794 1.0105 0.8892 0.3263 0.3894 1.0445 0.9454
WN_CONCEPT 0.1490 0.0242 0.2175 0.0522 0.3426 0.0218 0.2168 0.7791 1.0100 0.8836 0.3263 0.3894 1.0427 0.9390
ASJP SVD 0.5859 0.0253 1.0597 0.0522 1.0395 0.0225 1.3164 2.8254 2.9730 3.3461 0.0820 1.6568 3.8882 5.0063
ASJP UMAP 0.9318 0.0994 1.6767 0.1290 1.6055 0.0946 1.9856 3.3217 3.5364 4.0245 0.0811 2.3479 4.3690 5.5711
Östling and Tiedemann (2017)
L1 0.5697 0.0504 0.9843 0.0920 0.9616 0.0483 1.2819 4.7062 5.3973 5.9836 0.2837 1.7125 7.1117 7.7482
L2 0.5644 0.0497 0.9749 0.0918 0.9536 0.0452 1.2710 4.6788 5.4149 5.9684 0.1421 1.7019 7.1158 7.7296
L3 0.5589 0.0534 0.9601 0.0981 0.9336 0.0523 1.2558 4.7099 5.4686 6.0242 0.1138 1.6944 7.1689 7.7895
ALL 0.5595 0.0471 0.9673 0.0900 0.9437 0.0450 1.2641 4.6870 5.4175 5.9885 0.1482 1.6916 7.1240 7.7525
Table 15: KL Divergence between Language Similarity Graphs and the Language Confusion Matrices for Target/ Eval Languages from LCB and Inversion Tasks. The best results (lowest) are bolded, and the second best are underlined.